How to prepare 500mL of a 0.30M copper(II) sulfate solution ?
Answer #1

molarity = weight of substance*1000/G.m.Wt *valume of solution in ml

weight of substance           = molarity *G.M.Wt*valume of solution in ml/1000

                                            = 0.3*159.6*500/1000 = 23.94g

23.94 g of CuSO4 dissolve in 500ml of slution

0.15 moles of CuSO4 dissolve in 500ml of solution
Answer #2

number of moles of copper (II) sulfate solution is 0.3*0.5=0.15 mole

1 mole of copper (II) sulfate mass = 159.6gram

0.15 mole of copper (II) sulfate mass =0.15*159.6=23.94 grams